OpenStudy (anonymous):

How could I find the line passing through (2, -2) and (-5, 12)?

OpenStudy (anonymous):

y--2/12--2= x-2/-5-2 y+2/14= x-2/-7 cross multiplying -7(y+2)= 14( x-2) -7y -14= 14x-28 -7y-14x= -28+14 -7y-14x= -14 y+2x=2 2x+y=2

OpenStudy (anonymous):

Thanks